OpenStudy (anonymous):

The rate constant for a certain chemical reaction is 65.0 s-1 at a temperature of 20.0°C. If the activation energy for this reaction is 52. kJ/mole, what is the rate constant at a temperature of 30.0°C?

OpenStudy (aaronq):

\[\ln (\frac{ k _{2} }{ k _{1} })=\frac{ Ea }{ R }*\frac{ T _{2}-T_{1} }{ T _{2}*T _{1} }\]

OpenStudy (anonymous):

wat was the answer